OpenStudy (anonymous):

how would set up this equation? what mass of sucrose is needed to make 500ml of 2M solution? Thanks!

OpenStudy (frostbite):

Well we set up the equations we know and just substitute for unknown values :) \[\large C=\frac{ n }{ V }\] We know the concentration and volume, so we need to work with the amount of substance: \[\large n=\frac{ m }{ M }\] M is a constant so we get an expression with only 1 variable, the mass: \[\large CV=n \to CV = \frac{ m }{ M }\] \[\large CVM=m\]

OpenStudy (anonymous):

what is this formula called?

OpenStudy (frostbite):

You can call it stoichiometry if you like.

OpenStudy (frostbite):

Or quantity calculations as I like to call it.

OpenStudy (frostbite):

Just remember to convert the units so they fit, when doing the calculation. :)